OpenStudy (anonymous):

A card was selected at random from a standard deck of cards. The suit of the card was recorded, and then the card was put back in the deck. The table shows the results after 40 trials. What is the relative frequency of selecting a heart? Outcome Club Diamond Heart Spade Number of trials 8 12 14 6 A. 15% B. 25% C. 27% D. 35%

OpenStudy (anonymous):

@iGreen @Michele_Laino

OpenStudy (michele_laino):

hint: the requated probability is given by the subsequent ratio: \[\large probability = \frac{{favorable\;cases}}{{possible\;cases}}\]

OpenStudy (anonymous):

B?

OpenStudy (michele_laino):

oops.. requested*

OpenStudy (michele_laino):

no, since the favorable cases are 14, whereas the possible cases are: 14+12+8+6=40

OpenStudy (anonymous):

A?

OpenStudy (michele_laino):

\[\large probability = \frac{{favorable\;cases}}{{possible\;cases}} = \frac{{14}}{{40}} = ...?\]

OpenStudy (anonymous):

D?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

YAY can you help me with 4 more please?

OpenStudy (michele_laino):

ok!